OpenStudy (anonymous):

mixture tables- A candy mixture is to be made from chocolate costing $5 per pound and sugar candies costing $2 per pound. If a total mixture of 50 pounds of mixture is required at a cost of $3 per pound, how much of each must go into the mixture?

OpenStudy (anonymous):

Define your variables: let C=number pounds of chocolate let S=number of pounds of sugar candies Then the variables can be related according to total weight: C+S=50 they may also be related according to cost: 5C+2S=3(C+S), which reduces to 2C=S Sove the following system by substitution method C+S=50 [1] 2C=S [2] Substitute equation [2] into equation [1], and solve for C: C+2C=50 C=50/3 Substutute the value for C into [1] to find S: S=2(50/3)=100/3 Thus, to get 50 pounds of the candy blend at $3/lb, you would need 50/3 or 16 2/3 pounds of chocolate and 100/3 or 33 1/3 pounds of sugar candies.